Please help me solve this equation: Find the perimeter of a rectangle whose length is 5 cm less than twice it's width. It's width is 15 cm.
Answer by mananth(16946)

width =15
length = 2*15-5=25
Perimeter = 2(L+W)
Perimeter = 2(25+15)
=2(40)
= 80 cm